Find the total force on the side of a water-filled tube 2.34 cm high with a radius of 0.300 cm.
1 : 0.506 N
2 : 0.0506 N
3 : 3.41 N
4 : 1.03 N
1
Expert's answer
2020-06-22T11:11:50-0400
"dA=2\pi rdh"

"dF=\rho gh dA=2\pi r\rho ghdh"

"F=\pi rH^2\rho g\\=\pi (0.003)(0.0234)^2(1000)(9.8)=0.0506\ N"

Answer: 2 : 0.0506 N